KAGRA information

(1)About apparent binary range on the lock monitor

CAL group is now checking calibration data and model more precisely for binary range calculation.

For more confirmation, I and Yamamoto-kun decided to take a transfer function again as a very short observation break. The time will be decided by Yamamoto-kun.

(2)IFO status

ADS for ETMs during LSC_LOCK guardian and BPC during observation are working very well. The longest lock is now continuing over 4h 20m at present (09:38).

General information from KAGRA office

[Repeated information (same with yesterday)]

As for COVID-19,

If you work in the control room or on a shift, it was decided that everyone should wear a mask (last night, I discussed this with Prof. Ohashi and others).

We will be giving out a few masks to these people, so please come to the office. However, please be careful not to dispose of the product once, but to use it multiple times.

We will no longer have towels for common use from tomorrow. Paper towels are also in short supply, and it is difficult to buy them because they are out of stock. Please bring your own handkercheif.

If a state of emergency is declared, Kamioka staffs may also be forced to work from home.

Please take responsibility for the disposal of food, etc. in the refrigerator.

Thank you for your cooperation.

Observation Staus Report (linked with klog)

KAGRA/Observation/KAGRA-1/ObservationApril7toApril21/8th (last edited 2020-04-08 11:09:44 by miyoki)